OpenStudy (anonymous):

Consider the circle such that points A(-3,1) and B(-3,-5) are on the circle and the radius is 7 units. The center of the circle is somewhere in quadrant IV. A) What are the exact coordinates of the center?

hero (hero):

Begin with the standard form of a circle: \[(x - h)^2 + (y-k)^2 = r^2\] Input both points separately, creating a system of two equations: \[(-3-h)^2 + (1 - k)^2 = 7^2\] \[(-3 - h)^2 + (-5 - k)^2 = 7^2\]

hero (hero):

Solve for (h,k)